# Einstein metrics on homogeneous superspaces

### v1

@@ 4.4.4 Classification results
$[\,x_{1}:x_{2}:x_{3}\mid c\,]=\begin{cases}\big{\{}(\lambda g,\tfrac{c}{\lambda})\mid g=(x_{1},x_{2},x_{3});\,\lambda\in\mathbb{R}^{\times}\big{\}},\ &\text{if $x_{1}x_{2}x_{3}\neq 0$,}\\[4.26773pt] \emptyset,\ &\text{if $x_{1}x_{2}x_{3}=0$,}\end{cases}$
@@
$\mathcal{F}_{1}(u,v)$ =$=\begin{cases}\big{\{}\big{(}(x_{2}+x_{3},x_{2},x_{3}),0\big{)}\mid x_{2}+x_{3}\neq 0;\,x_{2}x_{3}\neq 0\big{\}},\ &\text{if $u=p\ \&\ v=q$,}\\[4.26773pt] \emptyset,\ &\text{otherwise,}\end{cases}$
$\mathcal{F}_{3}(u,v)$ =$=\begin{cases}\big{\{}\big{(}(x_{1},x_{2},x_{1}+x_{2}),0\big{)}\mid x_{1}+x_{2}\neq 0;\,x_{1}x_{2}\neq 0\big{\}},\ &\text{if $u=p\ \&\ v=q$,}\\[4.26773pt] \emptyset,\ &\text{otherwise.}\end{cases}$
@@
Case 1 ($\text{Case 1 }(p<q\leq m):$ $\qquad(g,c)\in\mathcal{C}_{1,1}\cup\mathcal{C}_{1,2}\cup\mathcal{C}_{1,3}\cup\mathcal{C}_{1,4}\cup\mathcal{F}_{3}(m-n,2(m-n)),$
Case 2 ($\text{Case 2 }(p<m<q):$ $\qquad(g,c)\in\mathcal{C}_{2,1}\cup\mathcal{C}_{2,2}\cup\mathcal{C}_{2,3}\cup\mathcal{C}_{2,4}\cup\mathcal{F}_{1}(n-m,2m)\cup\mathcal{F}_{3}(m-n,2n),$
Case 3 ($\text{Case 3 }(m\leq p<q):$ $\qquad(g,c)\in\mathcal{C}_{3,1}\cup\mathcal{C}_{3,2}\cup\mathcal{C}_{3,3}\cup\mathcal{C}_{3,4}\cup\mathcal{F}_{1}(3m-n,2m),$
@@
Every $\mathcal{C}_{k,l}$ in Theorem 4.11 is of the form $[\,x_{1}:x_{2}:x_{3}\mid c\,]$, so $\mathcal{C}_{k,l}=\emptyset$ if and only if $x_{1}x_{2}x_{3}=0$. For $q=2(m-n)$, for example, $\mathcal{C}_{1,3}=[\,m-n-p:-m+n+p:0\mid 1\,]=\emptyset$. Corollary 4.12 below summarises the result of analysing all possible linear relations between the parameters $m,n,p,q$.
@@
  • Two solutions up to scaling: $\mathcal{C}_{2,1},\mathcal{C}_{2,4},\quad$ if $\ p=3(m-n)\ \&\ q=2n$.
@@
  • Two solutions up to scaling: $\mathcal{C}_{2,3},\mathcal{C}_{2,4},\quad$ if $\ p=n-m\ \&\ q=4m-2n$.
@@ 4.4.5 Positive metrics
$\mathcal{C}_{k,l}^{\pm}=\big{\{}(\lambda g,\tfrac{c}{\lambda})\mid g=(x_{1}^{k,l},x_{2}^{k,l},x_{3}^{k,l});\,c=c^{k,l};\,\lambda\in\mathbb{R}^{\pm}\big{\}},$
@@
$\mathcal{C}_{1,3}^{-}=\big{\{}\big{(}\lambda(m-n-p,m-n+p-q,2m-2n-q),\tfrac{1}{\lambda}\big{)}\mid\lambda<0\big{\}}.$
@@
$\mathcal{F}_{1}^{+}=\big{\{}\big{(}(x_{2}+x_{3},x_{2},x_{3}),0\big{)}\mid x_{2},x_{3}>0\big{\}},\qquad\mathcal{F}_{3}^{+}=\big{\{}\big{(}(x_{1},x_{2},x_{1}+x_{2}),0\big{)}\mid x_{1},x_{2}>0\big{\}}.$
@@
  • Continuous family of Ricci-flat solutions: $\mathcal{F}_{3}^{+},\quad$ if $\ p=m-n\ \&\ q=2(m-n)$.
@@
  • Four solutions up to scaling: $\begin{cases}\mathcal{C}_{2,1}^{+},\mathcal{C}_{2,2}^{+},\mathcal{C}_{2,3}^{+},\mathcal{C}_{2,4}^{+},\quad\text{if }\ n-m<m+n-q<p<m-n,\\[4.26773pt] \mathcal{C}_{2,1}^{-},\mathcal{C}_{2,2}^{-},\mathcal{C}_{2,3}^{-},\mathcal{C}_{2,4}^{-},\quad\text{if }\ m-n<p<m+n-q<n-m.\end{cases}$
@@
  • Continuous family of Ricci-flat solutions: $\mathcal{F}_{1}^{+},\quad$ if $\ p=3m-n\ \&\ q=2m$.
@@ 4.5.1 Isotropy summands
$\mathfrak{m}_{k}^{\mathbb{C}}:=\bigoplus_{\alpha\in\Delta_{k}}\mathfrak{g}_{\alpha}^{\mathbb{C}},\qquad\Delta_{k}:=\big{\{}\sum_{i=1}^{n+1}c_{i}\alpha_{i}\in\Delta_{M}\mid c_{p}=k\big{\}},\qquad k\in I_{\mathfrak{m}^{\mathbb{C}}},$
@@
Let ±$\pm\varepsilon,\pm\delta_{1},\dots,\pm\delta_{n}$ denote the weights of $\nu_{2|2n}$. In accord with the regular embedding $\mathfrak{k}^{\mathbb{C}}\subseteq\mathfrak{g}^{\mathbb{C}}$, we denote by $\varepsilon,\delta_{1},\dots,\delta_{p-1}$ the weights of $\mu_{1|p-1}$, by $-\varepsilon,-\delta_{1},\dots,-\delta_{p-1}$ the weights of $\bar{\mu}_{1|p-1}$, and by ±$\pm\delta_{p},\dots,\pm\delta_{n}$ the weights of $\nu_{0|2(n+1-p)}$. We then have
@@
the weights of $\mathfrak{m}_{2}^{{}^{\prime}\mathbb{C}}$ are $\delta_{i}+\delta_{j}$ with $11\leq i\leq j\leq p-$ 1 and $\varepsilon+\delta_{i}$ with $11\leq i\leq p-$ 1. As all these weights are contained in $\Delta_{2}$, it follows that $\mathfrak{m}_{2}^{{}^{\prime}\mathbb{C}}\subseteq\mathfrak{m}_{2}^{\mathbb{C}}$. We similarly have $\mathfrak{m}_{-1}^{{}^{\prime}\mathbb{C}}\subseteq\mathfrak{m}_{-1}^{\mathbb{C}}$ and $\mathfrak{m}_{-2}^{{}^{\prime}\mathbb{C}}\subseteq\mathfrak{m}_{-2}^{\mathbb{C}}$, so
@@
$\{A_{\alpha},B_{\alpha}\mid\alpha\in\Delta_{i}\cap\Delta_{\bar{0}}^{+}\}\cup\{\sqrt{\imath}A_{\alpha},\sqrt{\imath}B_{\alpha}\mid\alpha\in\Delta_{i}\cap\Delta_{\bar{1}}^{+}\}$
@@ 4.5.2 Structure constants and Ricci curvature
where $\{h_{1}=E_{11}-E_{22},\,h_{i}=E_{i+1,i+1}-E_{n+1+i,n+1+i}\mid 2\leq i\leq n+1\}$ is a basis for the Cartan subalgebra $\mathfrak{h}^{\mathbb{C}}$ (cf. (4.10)). The right $Q$-dual basis elements are given by
@@
Proposition 4.17 implies that $[112]$ is the only possibly nonzero structure constant up to permutation. Using Proposition 3.23 and the symmetry of $[ijk]$, we have $[112]=[211]=d_{2}(b_{2}-2c_{2})$. The result now follows using $b_{2}=b$, Proposition 4.18 and Lemma 4.20. ∎
Concretely, $b-2c_{1}=-(p-1)$ and $b-2c_{2}=-2(n+1-p)$, and since $2\leq p\leq n$, it follows that $(b-2c_{1})(b-2c_{2})>0$, $[112]\leq 0$ and that $[112]=0$ if and only if $p=2$.
@@ 4.5.3 Classification of Einstein metrics
[…]
@@
$[\,x_{1}:x_{2}\mid c\,]:=\begin{cases}\big{\{}(\lambda g,\tfrac{c}{\lambda})\mid g=(x_{1},x_{2});\,\lambda\in\mathbb{R}^{\times}\big{\}},\ &\text{if $x_{1}x_{2}\neq 0$,}\\[4.26773pt] \emptyset,\ &\text{if $x_{1}x_{2}=0$.}\end{cases}$
@@
The various expressions above, including the ones in Theorem 4.23, shorten if $p$ is replaced by $q+1$. This would correspond to enumerating the nodes associated to $\alpha_{2},\ldots,\alpha_{n}$ in the $\operatorname{SOSp}(2|2n)$ Dynkin diagram in Section 4.2.2 from $1$ to $n-1$. For consistency with the enumeration in the $\operatorname{SU}(m|n)$ cases in Section 4.3 and Section 4.4, we have opted to use $p$ as indicated.

### v2

@@ 4.4.4 Classification results
$[\,x_{1}:x_{2}:x_{3}\mid c\,]=\begin{cases}\big\{(\lambda g,\tfrac{c}{\lambda})\,|\,g=(x_{1},x_{2},x_{3});\,\lambda\in\mathbb{R}^{\times}\big\},\ &\text{if $x_{1}x_{2}x_{3}\neq 0$,}\\[4.26773pt] \emptyset,\ &\text{if $x_{1}x_{2}x_{3}=0$,}\end{cases}$
@@
$\mathcal{F}_{1}(u,v)$ $=\begin{cases}\big\{\big((x_{2}+x_{3},x_{2},x_{3}),0\big)\,|\,x_{2}+x_{3}\neq 0;\,x_{2}x_{3}\neq 0\big\},\ &\text{if $u=p\ \&\ v=q$,}\\[4.26773pt] \emptyset,\ &\text{otherwise,}\end{cases}$
$\mathcal{F}_{3}(u,v)$ $=\begin{cases}\big\{\big((x_{1},x_{2},x_{1}+x_{2}),0\big)\,|\,x_{1}+x_{2}\neq 0;\,x_{1}x_{2}\neq 0\big\},\ &\text{if $u=p\ \&\ v=q$,}\\[4.26773pt] \emptyset,\ &\text{otherwise.}\end{cases}$
@@
$\text{Case 1 }(p<q\leq m):$ $\qquad(g,c)\in\mathcal{C}_{1,1}\cup\mathcal{C}_{1,2}\cup\mathcal{C}_{1,3}\cup\mathcal{C}_{1,4}\cup\mathcal{F}_{3}(m-n,2(m-n)),$
$\text{Case 2 }(p<m<q):$ $\qquad(g,c)\in\mathcal{C}_{2,1}\cup\mathcal{C}_{2,2}\cup\mathcal{C}_{2,3}\cup\mathcal{C}_{2,4}\cup\mathcal{F}_{1}(n-m,2m)\cup\mathcal{F}_{3}(m-n,2n),$
$\text{Case 3 }(m\leq p<q):$ $\qquad(g,c)\in\mathcal{C}_{3,1}\cup\mathcal{C}_{3,2}\cup\mathcal{C}_{3,3}\cup\mathcal{C}_{3,4}\cup\mathcal{F}_{1}(3m-n,2m),$
@@
Every $\mathcal{C}_{k,l}$ in Theorem 4.11 is of the form $x_{1}:x_{2}:x_{3}\mid c\,]$, so $\mathcal{C}_{k,l}=\emptyset$ if and only if $x_{1}x_{2}x_{3}=0$. For q=2(mn)q=2(m-n), for example, $\mathcal{C}_{1,3}=[\,m-n-p:-m+n+p:0\mid 1\,]=\emptyset$. Corollary 4.12 below summarises the result of analysing all possible linear relations between the parameters $m,n,p,q$.
@@
  • Two solutions up to scaling: $\mathcal{C}_{2,1},\mathcal{C}_{2,4},\quad$ if p=3(mn)&q=2n\ p=3(m-n)\ \&\ q=2n.
@@
  • Two solutions up to scaling: $\mathcal{C}_{2,3},\mathcal{C}_{2,4},\quad$ if p=nm&q=4m2n\ p=n-m\ \&\ q=4m-2n.
@@ 4.4.5 Positive metrics
$\mathcal{C}_{k,l}^{\pm}=\big\{(\lambda g,\tfrac{c}{\lambda})\,|\,g=(x_{1}^{k,l},x_{2}^{k,l},x_{3}^{k,l});\,c=c^{k,l};\,\lambda\in\mathbb{R}^{\pm}\big\},$
@@
$\mathcal{C}_{1,3}^{-}=\big\{\big(\lambda(m-n-p,m-n+p-q,2m-2n-q),\tfrac{1}{\lambda}\big)\,|\,\lambda<0\big\}.$
@@
$\mathcal{F}_{1}^{+}=\big\{\big((x_{2}+x_{3},x_{2},x_{3}),0\big)\,|\,x_{2},x_{3}>0\big\},\qquad\mathcal{F}_{3}^{+}=\big\{\big((x_{1},x_{2},x_{1}+x_{2}),0\big)\,|\,x_{1},x_{2}>0\big\}.$
@@
  • Continuous family of Ricci-flat solutions: $\mathcal{F}_{3}^{+},\quad$ if p=mn&q=2(mn)\ p=m-n\ \&\ q=2(m-n).
@@
[…]
@@
  • Continuous family of Ricci-flat solutions: $\mathcal{F}_{1}^{+},\quad$ if p=3mn&q=2m\ p=3m-n\ \&\ q=2m.
@@ 4.5.1 Isotropy summands
$\mathfrak{m}_{k}^{\mathbb{C}}:=\bigoplus_{\alpha\in\Delta_{k}}\mathfrak{g}_{\alpha}^{\mathbb{C}},\qquad\Delta_{k}:=\big\{\sum_{i=1}^{n+1}c_{i}\alpha_{i}\in\Delta_{M}\,|\,c_{p}=k\big\},\qquad k\in I_{\mathfrak{m}^{\mathbb{C}}},$
@@
Let $\pm\varepsilon,\pm\delta_{1},\dots,\pm\delta_{n}$ denote the weights of $\nu_{2|2n}$. In accord with the regular embedding $\mathfrak{k}^{\mathbb{C}}\subseteq\mathfrak{g}^{\mathbb{C}}$, we denote by $\varepsilon,\delta_{1},\dots,\delta_{p-1}$ the weights of $\mu_{1|p-1}$, by $\varepsilon,-\delta_{1},\dots,-\delta_{p-1}$ the weights of $\bar{\mu}_{1|p-1}$, and by $\pm\delta_{p},\dots,\pm\delta_{n}$ the weights of $\nu_{0|2(n+1-p)}$. We then have
@@
the weights of $\mathfrak{m}_{2}^{{}^{\prime}\mathbb{C}}$ are $\delta_{i}+\delta_{j}$ with $1\leq i\leq j\leq p-1$ and $\varepsilon+\delta_{i}$ with $1\leq i\leq p-1$. As all these weights are contained in $\Delta_{2}$, it follows that $\mathfrak{m}_{2}^{{}^{\prime}\mathbb{C}}\subseteq\mathfrak{m}_{2}^{\mathbb{C}}$. We similarly have $\mathfrak{m}_{-1}^{{}^{\prime}\mathbb{C}}\subseteq\mathfrak{m}_{-1}^{\mathbb{C}}$ and $\mathfrak{m}_{-2}^{{}^{\prime}\mathbb{C}}\subseteq\mathfrak{m}_{-2}^{\mathbb{C}}$, so
@@
$\{A_{\alpha},B_{\alpha}\,|\,\alpha\in\Delta_{i}\cap\Delta_{\bar{0}}^{+}\}\cup\{\sqrt{\imath}A_{\alpha},\sqrt{\imath}B_{\alpha}\,|\,\alpha\in\Delta_{i}\cap\Delta_{\bar{1}}^{+}\}$
@@ 4.5.2 Structure constants and Ricci curvature
where $\{h_{1}=E_{11}-E_{22},\,h_{i}=E_{i+1,i+1}-E_{n+1+i,n+1+i}\,|\,2\leq i\leq n+1\}$ is a basis for the Cartan subalgebra $\mathfrak{h}^{\mathbb{C}}$ (cf. (4.10)). The right $Q$-dual basis elements are given by
@@
Proposition 4.17 implies that $[112]$ is the only possibly nonzero structure constant up to permutation. Using Proposition 3.25 and the symmetry of $[ijk]$, we have $112]=[211]=d_{2}(b_{2}-2c_{2})$. The result now follows using $b_{2}=b$, Proposition 4.18 and Lemma 4.20. ∎
Concretely, $b-2c_{1}=-(p-1)$ and $b-2c_{2}=-2(n+1-p)$, and since $2\leq p\leq n$, it follows that $(b-2c_{1})(b-2c_{2})>0$, $112]\leq 0$ and that $[112]=0$ if and only if $p=2$.
@@ 4.5.3 Classification of Einstein metrics
$(\mathrm{S1}):\quad(x_{1},x_{2})=(1,2),\qquad\ \ c=c_{1},$
$(\mathrm{S2}):\quad(x_{1},x_{2})=(c_{1},c_{2}),\qquad c=1+\frac{(2c_{1}-c_{2})c_{2}}{4c_{1}^{2}}.$
@@
$[\,x_{1}:x_{2}\mid c\,]:=\begin{cases}\big\{(\lambda g,\tfrac{c}{\lambda})\,|\,g=(x_{1},x_{2});\,\lambda\in\mathbb{R}^{\times}\big\},\ &\text{if $x_{1}x_{2}\neq 0$,}\\[4.26773pt] \emptyset,\ &\text{if $x_{1}x_{2}=0$.}\end{cases}$
@@
The various expressions above, including the ones in Theorem 4.23, shorten if $p$ is replaced by $q+1$. This would correspond to enumerating the nodes associated to $\alpha_{2},\ldots,\alpha_{n}$ in the $\operatorname{SOSp}(2|2n)$ Dynkin diagram in Section 4.2.2 from $1$ to n1n-1. For consistency with the enumeration in the $\operatorname{SU}(m|n)$ cases in Section 4.3 and Section 4.4, we have opted to use $p$ as indicated.